# Nuclear matter properties from chiral-scale effective theory including a dilatonic scalar meson

### v1

@@ IV Conclusion and discussion
Without introducing many freedoms, such as the $\delta$ meson, and with operators organized to respect chiral and scale symmetry considerations and expanded by chiral-scale orders, the bsHLS can provide a reasonable behavior of NM properties from subsaturation to intermediate densities, e.g., $K(n)$ and $E_{\rm sym}(n)$, compared to Walecka-type models. The NS structures are sensitive to NM properties in these density regions, making bsHLS outperform Walecka-type models in describing a wider range of densities. More specifically, the kink behavior of the $\sigma$ field in bsHLS at intermediate densities allows the $M_{\rm max}$ to reach nearly $3\ M_{\odot}$, while other NS observational constraints are still statisfied.
Besides, the behaviors of symmetry patterns in dense environments are also found to be pivotal to macroscopic phenomena: If there is no restoration point of scale symmetry at certain densities, such as the $\beta^{\prime}$ value of bsHLS-L, the NS structures will fall outside observational constraints; The $M_{\rm max}$ of predicted NSs is influenced by the behavior of order parameter of scale symmetry, $\langle\chi\rangle^{*}$. Furthermore, the study on the flow of $\langle\chi\rangle^{*}$ with densities suggests the necessity of introducing an additional suppression factor for $\tilde{g}_{\omega NN}$ which could be an interesting problem for further investigation.
In summary, introducing bsHLS to NM studies is a promising approach due to its close relation to QCD symmetry patterns and the effective potentials organized by chiral-scale orders, which have already proven successful in describing scattering experiments at vacuum. Furthermore, the difference between bsHLS and Walecka-type models is not far from $n_{0}$, making it possible to be verified in future experiments, such as heavy ion collisions. The relationship between microscopic symmetries and macroscopic phenomena found in this work is also a valuable topic for further study.

### v3

@@ IV Conclusion and discussion
Without introducing many freedoms, such as the $\delta$ meson, and with operators organized to respect chiral and scale symmetry considerations and expanded by chiral-scale orders, the bsHLS can provide a reasonable behavior of NM properties from subsaturation to intermediate densities, e.g., $K(n)$ and $E_{\rm sym}(n)$, compared to Walecka-type models. And the NS structures are sensitive to NM properties at these density regions, making bsHLS outperform Walecka-type models in describing a wider range of densities. More specifically, the kink behavior of the $\sigma$ field in bsHLS at intermediate densities allows the $M_{\rm max}$ to reach nearly $3\ M_{\odot}$ for PNM, while other NS observational constraints are still statisfied.
Besides, the behaviors of symmetry patterns in dense environments are also found to be pivotal to macroscopic phenomena: If there is no restoration point of scale symmetry at certain densities, such as the $\beta^{\prime}$ value of bsHLS-L, the NS structures will fall outside observational constraints; The $M_{\rm max}$ of predicted NSs is influenced by the behavior of order parameter of scale symmetry, $\langle\chi\rangle^{*}$. Furthermore, the study on the flow of $\langle\chi\rangle^{*}$ with densities suggests the necessity of introducing an additional suppression factor for $\tilde{g}_{\omega NN}$, and it could be an interesting problem for further investigation.
In summary, introducing bsHLS to NM studies is a promising approach due to its close relation to QCD symmetry patterns and the effective potentials organized by chiral-scale orders, which have already proven successful in describing scattering experiments at vacuum. Furthermore, the difference between bsHLS and Walecka-type models is not far from $n_{0}$, making it possible to be verified in future experiments, such as heavy ion collisions. The relationship between microscopic symmetries and macroscopic phenomena found in this work is also a valuable topic to be further studied.
@@ Acknowledgments
We would like to thank Yi-Zhong Fan for his valuable comments.